# Geometry-Aware Decoding with Wasserstein-Regularized Truncation and Mass Penalties for Large Language Models

### v1

@@ 1 Introduction
Computing exact $W_{1}$ inside decoding is infeasible at vocabulary scale, so we optimize a tight dual-based surrogate that replaces the transport term with nearest-set distances $\mathrm{dist}(i,S)=\min_{j\in S}d(i,j)$ (Lemma 4.2). This yields simple, geometry-aware per-token scores and leads to an efficient alternating procedure: an $f$-step updates the potential (equivalently, the per-token geometric penalties/bonuses induced by the current crop), and an $S$-step updates the crop. For fixed $f$, we show the optimal crop is prefix-form after sorting tokens by the scalar score $i\mapsto f(i)+\lambda\log p_{i}$ (Theorem 3.4), so selecting $S$ reduces to scanning prefixes rather than searching over $2^{|\mathcal{V}|}$ subsets. Moreover, Top-$W$ provides a unifying view: by appropriate choices of the underlying metric, Top-$W$ reduces to familiar mass/entropy truncations such as Top-$k$ and Top-$H$ (Section 4.3). We evaluate Top-$W$ at fixed temperature on GSM8K, GPQA, AlpacaEval, and MT-Bench (plus LLM-as-a-judge creative writing) across 99 $(T,\text{model})$ settings with $T\in\{1.0,1.5,2.0\}$ and 3 LLMs. Across these settings, Top-$W$ consistently outperforms prior state-of-the-art decoding approaches: $9/9$ (GSM8K), $8/9$ (GPQA), $8/9$ (AlpacaEval), and $6/9$ (MT-Bench), delivering gains in both accuracy and judged creativity, while adding only modest runtime overhead (ms/token).
@@ 3 Top-$W$ Decoding
This section turns geometry-aware truncation into a computable next-token rule. In §3.1, we define the Wasserstein–entropy–mass objective $F_{\lambda,\beta}(S)$ and prove an exact factorization (Lemma 3.1), which makes the roles of geometry, entropy, and retained mass explicit. Optimizing $F_{\lambda,\beta}(S)$ is challenging because computing $W_{1}$ exactly would require solving the KR dual in 4 (or equivalently the OT linear program) at each decoding step, which is infeasible at vocabulary scale [22]. Instead, we propose a lightweight alternating scheme: an $f$-step that refines a feasible potential $f$, and an $S$-step that, for fixed $f$, optimizes a tight lower-bound surrogate in closed form. In §3.2 we focus on the $S$-step and show that for any fixed feasible $f$, minimizing the surrogate over $S$ is equivalent to maximizing a simple set function $G_{f}(S)$ (Lemma 3.2), which depends on $S$ only through its retained mass $\Gamma_{S}$. This yields a sharp structural characterization of the optimizer (Theorem 3.4): the optimal crop is obtained by scanning prefixes of tokens sorted by $\phi_{i}$, giving an $O(n)$ update rather than an $O(2^{n})$ search (over a vocabulary of size $n=|\mathcal{V}|$). Moreover, we show monotonicity of the selected retained mass in $\beta$ and $\lambda$ (Corollaries 3.5LABEL:cor:lambda-monotone-main), which directly motivates our logits-processor implementation.
@@ 3.1 Wasserstein–Entropy–Mass Objective and an Exact Factorization
Let $S\subseteq V$ with $\Gamma_{S}\in(0,1)$. Then
@@
For proof of Lemma 3.1, see Appendix B. By Lemma 3.1 and the entropy identity (3), minimizing (5) is equivalent to minimizing the expanded form
@@ 4.4 Mass parameter $\beta$: practical guidance
Table 1: GSM8K accuracy (%) at $T\in\{1.0,1.5,2.0\}$ for Min-$p$, Top-$p$, Top-$H$, and Top-$W$. Aggregated over 3 runs (N=1319).
@@
Table 2: GPQA accuracy at $T\in\{1.0,1.5,2.0\}$ for Min-$p$, Top-$p$, Top-$H$, and Top-$W$. Aggregated over 4 runs $(N=448)$.
@@ 5 Experimental Details
We evaluate on both accuracy-centric and judge-based benchmarks: GSM8K [5] and GPQA [25] are scored by exact-match accuracy (%) from extracted final answers (GSM8K) or the chosen option (GPQA). For AlpacaEval, we report length-controlled win rates using the standard length-debiasing protocol [9]. For MT-Bench, we report the average judge score (1–10) following the canonical pipeline [33]. To probe quality axes beyond task accuracy, we also run a rubric-based LLM-as-Judge evaluation on three creative prompts: for each prompt and temperature, we generate one response per method, score with a single judge across five rubrics (M1–M5), aggregate to an overall score, and randomize method order to mitigate positional bias. Appendix L.1 contains the prompt templates and rubrics.
@@
All methods are evaluated at temperatures $T\in\{1.0,1.5,2.0\}$ with identical max-token budgets, prompts/format constraints, and answer-extraction rules; only the truncation/cropping rule differs. We run GSM8K and GPQA via lm-eval-harness with benchmark-standard prompting and extraction [11]. For AlpacaEval, MT-Bench, and the rubric-based creative evaluation, we use GPT-4o as the judge and randomize candidate order per item (and per repeat, when applicable). More details including judge prompts is in Appendix L.1.
@@
Top-$W$ applies cropping to the temperature-scaled next-token distribution and is controlled by $\lambda$ (entropy penalty) and $\beta$ (log-mass weight). Unless stated otherwise, we fix $\lambda=2.2$ and $\beta=2.8$ (see §6.2 and Appendix M). In implementation, we use a nucleus-style warm start, a fixed top_m=1200=1200 candidate pool, and alt_iters=3=3 alternating refinement steps; these settings are fixed across tasks unless stated otherwise.
@@ 6 Experiments
We evaluate whether Top-$W$ consistently improves over Min-$p$, Top-$p$ and Top-$H$ under matched temperature, prompts, max-token budgets, and stopping criteria. We report exact-match accuracy on GSM8K and GPQA, and judge-based outcomes on AlpacaEval, MT-Bench, and a multi-axis creative-writing rubric. We additionally sweep $(\lambda,\beta)$ to probe sensitivity and the accuracy–diversity trade-off. For GSM8K and GPQA, we quote the Min-$p$, Top-$p$ and Top-$H$ accuracy results from Potraghloo et al. [23]. For AlpacaEval and MT-Bench, absolute scores depend on the judge model/prompt; we therefore run Min-$p$, Top-$p$, Top-$H$ (using their provided implementation), and Top-$W$ under our fixed judge configuration and protocol (Appendix J). We use the authors’ released implementation of Top-$H$ and reproduce Min-pp/Top-$p$ under the same evaluation pipeline; for GSM8K and GPQA we additionally quote baseline numbers reported in Potraghloo et al. [23].
Figure 1: Alpaca accuracy across temperatures for Min-$p$, Top-$p$, Top-$H$, and Top-$W$ (aggregated over 4 runs). As we can see in the bar plot, Min-$p$, Top-$p$, Top-$H$ and Top-$W$ (our method) win in 0,0,1,8 tuples of $(T,model)$ out of 9 tuples, respectively.
Figure 2: MT-Bench judge scores across temperatures for Min-$p$, Top-$p$, Top-$H$, and Top-$W$ (aggregated over 4 runs). As we can see in the bar plot, Min-$p$, Top-$p$, Top-$H$ and Top-$W$ (our method) win in 0,1,2,6 tuples of $(T,model)$ out of 9 tuples, respectively.
@@ 6.1 Main Results
Tables 1 and 2 report exact-match accuracy across $T\in\{1.0,1.5,2.0\}$. Across models, Min-$p$ and Top-$p$ degrade sharply as $T$ increases, while Top-$H$ and Top-$W$ are substantially more stable. On GSM8K, Top-$W$ beats Min-$p$, Top-$p$, and Top-$H$ across all three models (Qwen2.5-3B, LLaMA-3.1-8B, Phi-3-Mini) and all three temperatures $T\in\{1.0,1.5,2.0\}$; the improvement over Top-$H$ reaches up to $33.74\%$ at T=2.0T{=}2.0. Across all $T$, LLaMA-3.1-8B is the strongest absolute performer, while Qwen2.5-3B and Phi-3-Mini benefit more from Top-$W$’s stability at higher $T$. GSM8K (multi-step arithmetic) is more sensitive to sampling noise, so probability-only truncators collapse faster than on GPQA, whereas Top-WW/Top-$H$ preserve accuracy as $T$ rises. On GPQA (knowledge-heavy multiple-choice), accuracies are lower and tighter across methods, but Top-$W$ delivers the most consistent gains, winning $8/9$ $(T,\text{model})$ settings: it outperforms baselines in Qwen2.5-3B and Phi-3-Mini for all $T$, and improves over LLaMA-3.1-8B at $T\in\{1.5,2.0\}$ while trailing Top-$p$ at T=1.0T{=}1.0.
@@
Figures 12 summarize judge-based performance across temperatures and models. Top-$W$ achieves the strongest overall scores and remains robust as $T$ increases; it also wins the majority of $(T,\text{model})$ tuples (e.g., $8/9$ on AlpacaEval and $6/9$ on MT-Bench in our reported aggregates). We observe larger method gaps on MT-Bench than AlpacaEval for some models, consistent with MT-Bench’s multi-turn coherence and instruction retention being more brittle under increased stochasticity. AlpacaEval often rewards single-turn helpfulness/style, while MT-Bench more strongly penalizes drift and inconsistency, which amplifies the value of Top-$W$’s geometry-aware truncation at higher $T$.
@@ 6.2 Top-$W$ Ablations and Behavior
Focusing on LLaMA-3.1-8B (because of its more divers behavior on Top-$W$) and GSM8K, we run a parameter sensitivity study over $\beta$ (and $\lambda$) to quantify how Top-$W$ transitions from conservative, accuracy-favoring behavior to more diverse generations (Figure 3). In particular, large $\beta$ can increase diversity in creative settings (as reflected by rubric-based judging), while potentially degrading performance on strict-answer reasoning tasks. As summarized by the sweep plots and the rubric-based judge tables (Tables 8 and 3), $\beta$ acts as a control knob on the exploration–conservatism axis: smaller $\beta$ tends to favor sharper truncation and higher accuracy on strict-answer tasks, while larger $\beta$ expands the viable support and can improve diversity- and style-sensitive judgments.
@@ Appendix I Candidate-pool exactness theorem (justifying Top-M restriction)
Appendix L Results of LLM as a judge
L.1 LLM-as-a-Judge evaluation prompts
@@
L.2 Results
Below, we go over per-model breakdown of Table 8. Judge-based evaluation metrics (scores from 1.0 to 10.0) across temperatures, prompts, and sampling methods are reported for (Table 5) LLaMA3.1-8B-Instruct, (Table 6) Phi-3-Mini and (Table 7) Qwen2.5-3B. M1–M5 correspond to creativity, originality, narrative flow, imagery, and vitality, respectively.

### v2

@@ 1 Introduction
Computing exact $W_{1}$ inside decoding is infeasible at vocabulary scale, so we optimize a tight dual-based surrogate that replaces the transport term with nearest-set distances $\mathrm{dist}(i,S)=\min_{j\in S}d(i,j)$ (Lemma 4.2). This yields simple, geometry-aware per-token scores and leads to an efficient alternating procedure: an $f$-step updates the potential (equivalently, the per-token geometric penalties/bonuses induced by the current crop), and an $S$-step updates the crop. For fixed $f$, we show the optimal crop is prefix-form after sorting tokens by the scalar score $i\mapsto f(i)+\lambda\log p_{i}$ (Theorem 3.4), so selecting $S$ reduces to scanning prefixes rather than searching over $2^{|\mathcal{V}|}$ subsets. Moreover, Top-$W$ provides a unifying view: by appropriate choices of the underlying metric, Top-$W$ reduces to familiar mass/entropy truncations such as Top-$k$ and Top-$H$ (Section 4.3). We evaluate Top-$W$ at fixed temperature on GSM8K, GPQA, AlpacaEval, and MT-Bench (plus LLM-as-a-judge creative writing) across 1515 $(T,\text{model})$ settings with $T\in\{0.5,0.7,1.0,1.5,2.0\}$ and 3 LLMs. Across these settings, Top-$W$ consistently outperforms prior state-of-the-art decoding approaches: $13/15$ (GSM8K), $12/15$ (GPQA), $12/15$ (AlpacaEval), and $8/15$ (MT-Bench), delivering gains in both accuracy and judged creativity, while adding only modest runtime overhead (ms/token).
@@ 3 Top-$W$ Decoding
This section turns geometry-aware truncation into a computable next-token rule. In §3.1, we define the Wasserstein–entropy–mass objective $F_{\lambda,\beta}(S)$ and prove an exact factorization (Lemma 3.1), which makes the roles of geometry, entropy, and retained mass explicit. Optimizing $F_{\lambda,\beta}(S)$ is challenging because computing $W_{1}$ exactly would require solving the KR dual in 4 (or equivalently the OT linear program) at each decoding step, which is infeasible at vocabulary scale [24]. Instead, we propose a lightweight alternating scheme: an $f$-step that refines a feasible potential $f$, and an $S$-step that, for fixed $f$, optimizes a tight lower-bound surrogate in closed form. In §3.2 we focus on the $S$-step and show that for any fixed feasible $f$, minimizing the surrogate over $S$ is equivalent to maximizing a simple set function $G_{f}(S)$ (Lemma 3.2), which depends on $S$ only through its retained mass $\Gamma_{S}$. This yields a sharp structural characterization of the optimizer (Theorem 3.4): the optimal crop is obtained by scanning prefixes of tokens sorted by $\phi_{i}$, giving an $O(n)$ update rather than an $O(2^{n})$ search (over a vocabulary of size $n=|\mathcal{V}|$). Moreover, we show monotonicity of the selected retained mass in $\beta$ (Corollary 3.5), which directly motivates our logits-processor implementation.
@@ 3.1 Wasserstein–Entropy–Mass Objective and an Exact Factorization
Let $S\subseteq V$ with $\Gamma_{S}\in(0,1)$ and define $S^{c}$ as its complement in $V$. Then
@@
$W_{1}(p,q_{S})$ separates into (i) the amount of removed probability mass, $(1-\Gamma_{S})$ , and (ii) the geometric distance between the removed and retained conditional distributions. For proof of Lemma 3.1, see Appendix B. By Lemma 3.1 and the entropy identity (3), minimizing (5) is equivalent to minimizing the expanded form
@@ 4.4 Mass parameter $\beta$: practical guidance
Table 1: GSM8K accuracy (%) at $T\in\{0.5,0.7,1.0,1.5,2.0\}$ for Min-$p$, Top-$p$, Top-$H$, and Top-$W$. Aggregated over 3 runs (N=1319).
@@
Table 2: GPQA accuracy at $T\in\{0.5,0.7,1.0,1.5,2.0\}$ for Min-$p$, Top-$p$, Top-$H$, and Top-$W$. Aggregated over 4 runs $(N=448)$.
@@ 5 Experimental Details
We evaluate on both accuracy-centric and judge-based benchmarks: GSM8K [6] and GPQA [26] are scored by exact-match accuracy (%) from extracted final answers (GSM8K) or the chosen option (GPQA). For AlpacaEval, we report length-controlled win rates using the standard length-debiasing protocol [10]. For MT-Bench, we report the average judge score (1–10) following the canonical pipeline [32]. To probe quality axes beyond task accuracy, we also run a rubric-based LLM-as-Judge evaluation on three creative prompts: for each prompt and temperature, we generate one response per method, score with a single judge across five rubrics (M1–M5), aggregate to an overall score, and randomize method order to mitigate positional bias. Appendix M.1 contains the prompt templates and rubrics.
@@
All methods are evaluated at temperatures $T\in\{0.5,0.7,1.0,1.5,2.0\}$ with identical max-token budgets, prompts/format constraints, and answer-extraction rules; only the truncation/cropping rule differs. We run GSM8K and GPQA via lm-eval-harness with benchmark-standard prompting and extraction [12]. For AlpacaEval, MT-Bench, and the rubric-based creative evaluation, we use GPT-4o as the judge and randomize candidate order per item (and per repeat, when applicable). More details including judge prompts is in Appendix M.1.
@@
Top-$W$ applies cropping to the temperature-scaled next-token distribution and is controlled by $\lambda$ (entropy penalty) and $\beta$ (log-mass weight). Unless stated otherwise, we fix $\lambda=2.2$ and $\beta=2.8$ (see §6.2 and Appendix L). In implementation, we use a nucleus-style warm start, a fixed top_m=1200=1200 candidate pool, and alt_iters=3=3 alternating refinement steps; these settings are fixed across tasks unless stated otherwise.
@@ 6 Experiments
We evaluate whether Top-$W$ consistently improves over Min-$p$, Top-$p$ and Top-$H$ under matched temperature, prompts, max-token budgets, and stopping criteria. We report exact-match accuracy on GSM8K and GPQA, and judge-based outcomes on AlpacaEval, MT-Bench, and a multi-axis creative-writing rubric. We additionally sweep $(\lambda,\beta)$ to probe sensitivity and the accuracy–diversity trade-off. For GSM8K and GPQA, we quote the Min-$p$, Top-$p$ and Top-$H$ accuracy results from Baghaei Potraghloo et al. [2]. For AlpacaEval and MT-Bench, absolute scores depend on the judge model/prompt; we therefore run Min-$p$, Top-$p$, Top-$H$ (using their provided implementation), and Top-$W$ under our fixed judge configuration and protocol (Appendix J). We use the authors’ released implementation of Top-$H$ and reproduce Min-pp/Top-$p$ under the same evaluation pipeline; for GSM8K and GPQA we additionally quote baseline numbers reported in Baghaei Potraghloo et al. [2].
Figure 1: Alpaca accuracy across temperatures for Min-$p$, Top-$p$, Top-$H$, and Top-$W$ (aggregated over 4 runs). As we can see in the bar plot, Min-$p$, Top-$p$, Top-$H$ and Top-$W$ (our method) win in 0,1,2,12 tuples of $(T,model)$ out of 15 tuples, i.e., 5 temperatures $\times$ 3 models, respectively.
Figure 2: MT-Bench judge scores across temperatures for Min-$p$, Top-$p$, Top-$H$, and Top-$W$ (aggregated over 4 runs). As we can see in the bar plot, Min-$p$, Top-$p$, Top-$H$ and Top-$W$ (our method) win in 1,4,2,8 tuples of $(T,model)$ out of 15 tuples, i.e., 5 temperatures $\times$ 3 models, respectively.
@@ 6.1 Main Results
Tables 1 and 2 report exact-match accuracy across $T\in\{0.5,0.7,1.0,1.5,2.0\}$. Across models, Min-$p$ and Top-$p$ degrade sharply as $T$ increases, while Top-$H$ and Top-$W$ are substantially more stable. On GSM8K, Top-$W$ wins (or mutually wins) $13/15$ $(T,\text{model})$ settings across three models (Qwen2.5-3B, LLaMA-3.1-8B, Phi-3-Mini) and five temperatures $T\in\{0.5,0.7,1.0,1.5,2.0\}$; the improvement over Top-$H$ reaches up to $33.74\%$ at T=2.0T{=}2.0. Across all $T$, LLaMA-3.1-8B is the strongest absolute performer, while Qwen2.5-3B and Phi-3-Mini benefit more from Top-$W$’s stability at higher $T$. GSM8K (multi-step arithmetic) is more sensitive to sampling noise, so probability-only truncators collapse faster than on GPQA, whereas Top-WW/Top-$H$ preserve accuracy as $T$ rises. On GPQA (knowledge-heavy multiple-choice), accuracies are lower and tighter across methods, but Top-$W$ still delivers the most consistent gains, winning (or mutually winning) $12/15$ $(T,\text{model})$ settings: it dominates for all three models at $T\in\{1.5,2.0\}$ and remains top or tied at T=1.0T{=}1.0, while margins shrink at $T\in\{0.5,0.7\}$ where baselines such as Min-$p$ (LLaMA-3.1-8B, T=0.5T{=}0.5) and Top-$H$ (LLaMA-3.1-8B, T=0.7T{=}0.7) occasionally win.
@@
Figures 12 summarize judge-based performance across temperatures and models. Top-$W$ achieves the strongest overall scores and remains robust as $T$ increases; it also wins the majority of $(T,\text{model})$ tuples (e.g., $12/15$ on AlpacaEval and $8/15$ on MT-Bench in our reported aggregates). We observe larger method gaps on MT-Bench than AlpacaEval for some models, consistent with MT-Bench’s multi-turn coherence and instruction retention being more brittle under increased stochasticity. AlpacaEval often rewards single-turn helpfulness/style, while MT-Bench more strongly penalizes drift and inconsistency, which amplifies the value of Top-$W$’s geometry-aware truncation at higher $T$.
@@ 6.2 Top-$W$ Ablations and Behavior
Focusing on LLaMA-3.1-8B (because of its more divers behavior on Top-$W$) and GSM8K, we run a parameter sensitivity study over $\beta$ (and $\lambda$) to quantify how Top-$W$ transitions from conservative, accuracy-favoring behavior to more diverse generations (Figure 3). In particular, large $\beta$ can increase diversity in creative settings (as reflected by rubric-based judging), while potentially degrading performance on strict-answer reasoning tasks. As summarized by the sweep plots and the rubric-based judge tables (Tables  3 and 8), $\beta$ acts as a control knob on the exploration–conservatism axis: smaller $\beta$ tends to favor sharper truncation and higher accuracy on strict-answer tasks, while larger $\beta$ expands the viable support and can improve diversity- and style-sensitive judgments.
@@ Appendix I Candidate-pool exactness theorem (justifying Top-M restriction)
Appendix L Hyperparameter Tuning
[…]
Appendix M Results of LLM as a judge
M.1 LLM-as-a-Judge evaluation prompts
@@
M.2 Results
Below, we go over per-model breakdown of Table 8. Judge-based evaluation metrics (scores from 1.0 to 10.0) across temperatures, prompts, and sampling methods are reported for (Table 5) LLaMA3.1-8B-Instruct, (Table 6) Phi-3-Mini and (Table 7) Qwen2.5-3B.